# A Benchmark for the Application of Distributed Control Techniques to the Electricity Network of the European Economic Area

### v1

@@ 4.1 Links to Sources, Limitations, Costs, and Licensing
The benchmark is implemented in Matlab (r2023b), and the necessary files to execute it are available at: https://gitlab.tudelft.nl/ariccardi/european-economic-area-electricity-network. Gurobi Optimizer [^1] https://www.gurobi.com/ is required for the computation of the centralized MPC strategy. Alternatively, the Matlab Optimization Toolbox [^2] https://www.mathworks.com/products/optimization can be used with minimal modifications.
@@ 4.2 Documentation
Documentation for the benchmark is available at: https://gitlab.tudelft.nl/ariccardi/ european-economic-area-electricity-network.
For clarity, in the following, we provide the user with information about the functions used. The data from _ENTSOE has been preliminarily checked for integrity, replacing missing entries with linear interpolations, and reported on a consistent scale. This process was performed with specialized scripts reported in the online documentation. The resulting preprocessed data set is also part of the benchmark and provided as .csv files. The benchmark is constituted by the following files:
@@ 5.1 Reference Approach: Centralized Predictive Control
were $\bar{\mathbf{u}}=\left[\mathbf{u}^{\intercal}(0|k)\cdots\mathbf{u}^{\intercal}(N-1|k)\right]^{\intercal}$, $\mathbf{R}$ and $\mathbf{Q}$ are diagonal cost matrices of appropriate dimensions, and $N$ is the prediction horizon. According to the receding horizon logic of MPC, we apply $\mathbf{u}(0|k)$ to the system, discard the remaining control actions, and iterate. Problem (17) is a quadratic optimization problem, hence efficient algorithms for solving it exist in the literature. For example, the problem can be solved using an active-set or an interior point algorithm. For the benchmark, the optimization is performed with Gurobi Optimizer using the barrier algorithm. The simulation required 206 [h], 15[m], and 24[s] to be completed using a processor Intel Xeon E5-2637v3, with a base clock of 3.5 GHz, and coupled with 128 GB of RAM. The solution of the optimal control problem is the vector of inputs $\mathbf{u}(k)$, for $k=0,...,34559$, (corresponding to 24 hours of simulation) reported in Figure 5.
Refer to caption (a) Input $u_{1}$. Refer to caption (b) Input $u_{3}$. Refer to caption (c) Input $u_{2}$. Refer to caption (d) Power transmitted over the tie lines.
@@
The evolution of the states resulting from the sequence of inputs obtained through MPC is reported in Figure 6. For the overall power balance of each country, please consult the online repository https://gitlab.tudelft.nl/ariccardi/european-economic-area-electricity-network.
The evolution of the cost function is presented in Figure 7. Its value is the metric to evaluate distributed control techniques w.r.t. the proposed centralized approach. In the application of distributed control, the performance of the system usually decreases to achieve auxiliary objectives such as improvement in the computation time required to obtain the control action, reduction in the volume of information exchanged across the network (with advantages in the sector of security too), or reliability and redundancy of control in the presence of faults or unforeseen events.
Refer to caption (a) State $x_{1}$. Refer to caption (b) State $x_{2}$. Refer to caption (c) State $x_{3}$. Refer to caption (d) State $x_{4}$.
@@ Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101018826) – Project CLariNet

### v3

@@ 4.1 Links to Sources, Limitations, Costs, and Licensing
The benchmark is implemented in Matlab (r2023b), and the necessary files to execute it are available at riccardi_CodeUnderlyingPublication_2024b . Gurobi Optimizer [^1] https://www.gurobi.com is required for the computation of the centralized MPC strategy. Alternatively, the Matlab Optimization Toolbox [^2] https://www.mathworks.com/products/optimization can be used with minimal modifications.
@@ 4.2 Documentation
Documentation for the benchmark is available at riccardi_CodeUnderlyingPublication_2024b . In the following, we provide the user with information about the functions used. The data from _ENTSOE has been preliminarily checked for integrity, replacing missing entries with linear interpolations, and reported on a consistent scale. This process was performed with specialized scripts reported in the online documentation. The resulting preprocessed data set is also part of the benchmark and provided as .csv files. The benchmark is constituted by the following files:
@@ 5.1 Reference Approach: Centralized Predictive Control
were $\bar{\mathbf{u}}=\left[\mathbf{u}^{\intercal}(0|k)\cdots\mathbf{u}^{\intercal}(N-1|k)\right]^{\intercal}$, $\mathbf{R}$ and $\mathbf{Q}$ are diagonal cost matrices of appropriate dimensions such that, for each area $i$, $\mathbf{R}_{i}=\text{diag}[100;\,10;\,1]$ and $\mathbf{Q}_{i}=\text{diag}[1;\,1;\,1]$, and $N=30$ is the prediction horizon. For the external signals $\mathbf{w}$ we use the current measurement for $j=1$, and their forecast for the remaining time steps. According to the receding horizon logic of MPC, we apply $\mathbf{u}(0|k)$ to the system, discard the remaining control actions, and iterate. Problem (12) is a quadratic optimization problem, hence efficient algorithms for solving it exist in the literature. For example, the problem can be solved using an active-set or an interior point algorithm.
For the benchmark, the optimization is performed with Gurobi Optimizer using the barrier algorithm. The simulation required 206 [h], 15[m], and 24[s]333 The amount of time required to run the simulation is only indicative, and serves to understand the complexity of the system. The end-user should re-run the benchmark on their own system to obtain results that are comparable with the architectures at study. to be completed using a processor Intel Xeon E5-2637v3, with a base clock of 3.5 GHz, and coupled with 128 GB of RAM. The solution of the optimal control problem is the vector of inputs $\mathbf{u}(k)$, for $k=0,...,34559$, (corresponding to 24 hours of simulation) reported in Fig. 5.
(a) Input $u_{1}$.
(b) Input $u_{3}$.
(c) Input $u_{2}$.
(d) Power transmitted over the tie lines.
@@
The evolution of the states resulting from the sequence of inputs obtained through MPC is reported in Fig. 6. For the overall power balance of each country, please consult the online repository riccardi_CodeUnderlyingPublication_2024b .
The evolution of the cost function is presented in Fig. 7. Its value is the metric to evaluate distributed control techniques w.r.t. the proposed centralized approach. In the application of distributed control, the performance of the system usually decreases to achieve auxiliary objectives such as improvement in the computation time required to obtain the control action, reduction in the volume of information exchanged across the network (with advantages in the sector of security too), or reliability and redundancy of control in the presence of faults or unforeseen events.
(a) State $x_{1}$.
(b) State $x_{2}$.
(c) State $x_{3}$.
(d) State $x_{4}$.
@@ Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101018826) – Project CLariNet